what is the force of attraction between a - 3kg mass and a 6.2kg object when they are seperated at 0.6km?
1
Expert's answer
2020-04-20T10:42:10-0400

The Newton's law of gravitation says

"F=G\frac{m_1m_2}{r^2}\\
=6.67\times 10^{-11}\frac{3.0\times 6.2}{(600)^2}=3.45\times 10^{-15}\:\rm N."